OpenStudy (anonymous):

What is the equation of a line that passes through the point (3, -1) and is perpendicular to a line with a slope of 3/4

OpenStudy (anonymous):

What is the equation of a line that passes through the point (0, -3) and is perpendicular to a line with a slope of 4?

OpenStudy (anonymous):

One question at a time. You want to put into the equation of a line form, which is y = mx+b. You are given x and y (3, -1) and slope (m) = 3/4. Just substitute these values into the equation to solve for b. Do this now. Once you solve for b, you can put everything together.

OpenStudy (anonymous):

sorry that was the wrong question What is the equation of a line that passes through the point (0, -3) and is perpendicular to a line with a slope of 4?

OpenStudy (anonymous):

Doesn't matter, it's the same process.

hero (hero):

@iambatman, the slope of the equation of the line is different from the slope of the perpendicular line.

OpenStudy (anonymous):

-3=4(-1)+b

OpenStudy (anonymous):

Right, just take the negative reciprocal for the slope.

OpenStudy (anonymous):

y=1/4+3